what adds to make -11 and times to make positive 10
Answer by JoelSchwartz(130)


x+y=-11
xy=10
y(-11-y)=10
-11y-y^2=10
-y^2-11y-10=0
ac=10
(y+10)(-y-1)=0
y=-10 and y =-1
-10-1=-11
-10*-1=10